# ADD 2023: Towards Audio Deepfake Detection and Analysis in the Wild

### v1

@@ V-C Track 2: Manipulation region location (RL)
    Acoustic features: Like in the FG-D task, most teams opted to use Wav2Vec 2.0 for acoustic feature extraction, with the notable exception of Teams C02 [43] and C03 [44], who used spectrograms as acoustic feature representation. This popularity of Wav2Vec 2.0 further evidences the reliability of deep embedding features in anti-deepfake systems. In addition, it is worth noting that Team C05 [46] uses both raw waveform as well as wav2vec deep embedding features.
@@
    Back-end classifiers: Given the sequence nature of the RL task, where a real/fake prediction must be given for each frame, LSTM-based models are popular among the top-performing teams, with the notable exception of Team C05 [46], who used an AASIST-based model. Judging by the ranking, however, it is evident that LSTM-based recurrent models may still be more effective and better suited for the recurrent nature of the RL task, as Team C05, who used a fusion of AASIST-based model and FC layer classifier, is the only team in the top-5 that did not use an LSTM-based model, and the performance is slightly lower than that of the other teams.
V-D Track 3: Deepfake algorithm recognition (AR)
@@
    Data augmentation: Most teams opted to use MUSAN for noise augmentation, and a significant number of teams also opted to add reverberation to their training data to improve the robustness of their models. In addition, the top-performing teams also opted to use other methods to further augment their training data, including CutMix [58] and random sampling, time stretching, and time masking. This is likely due to the fact that the AR task is concerned with the recognition of deepfake algorithms, and thus the use of different data augmentation methods may be more effective in capturing the diverse acoustic conditions of deepfake audio in the wild.
@@
    Acoustic features: Compared to Tracks 1.2 and 2, participants of Track 3 are more diverse in their choice of acoustic features. While a number of teams used Wav2Vec 2.0, a significant number of teams also opted to use log-Mel spectrogram, STFT spectrogram, and even HuBERT-based features. This is likely due to the fact that the AR task is concerned with the recognition of deepfake algorithms, beyond the binary classification of real/fake, and thus a more diverse set of acoustic features may be more suitable for the task, without a clear-cut winner. This also signals the prospect of additional research into the development of acoustic features for anti-deepfake systems.
@@
    Back-end classifiers: ResNet-based [59] models (including Res2Net [60]) are popular among the top-performing teams, with the notable exception of Team D03 [49], who used a single ECAPA-TDNN-based model [61]. The use of ECAPA-TDNN is likely due to the fact that ECAPA-TDNN is a popular model for speaker recognition, and the philosophy of speaker recognition may be applicable to the AR task, which is also a recognition task, albeit of deepfake algorithms instead of speakers. The popularity of ResNet-based models is likely due to the fact that ResNet-based models are relatively simple and efficient, and thus may be more suitable for real-world applications.
@@
    Model fusion: The averaging of scores is still popular, but the use of label fusion and feature fusion (in the form of concatenation) also see its use. This is likely due to the fact that the AR task is concerned with the recognition of deepfake algorithms, and thus the use of different models, each of which may be better suited for a particular aspect of the task, may be more effective than a single-model approach.
@@
    OSR methods: Given the nature of the AR task and its practicality in real-world applications, as well as the presence of an unknown class, participating teams are required to use an out-of-distribution detection method to detect unknown classes. Thresholding remains the most popular method, to decent success. Other methods, including OpenMax [62] and manifold-based multi-model fusion, are also used.
VI Future Directions
The ADD 2023 challenge has provided a platform for researchers to develop new technologies to combat deepfake audio; However, there are several limitations that should be addressed in future research efforts. We identify these limitations and suggest future directions for research in the field of deepfake audio detection.
  • Coping with unseen deepfake technologies and adversarial attacks: The rapid development of audio deepfake generation and adversarial attack technologies, like VALL-E, GPT-4o, VISinger and DiffSinger, etc., brings critical challenges to current existing detection methods. In response to these challenges, deepfake audio generation and detection task are viewed as a rivalry game for participants in the ADD competitions. Despite partly improving the anti-attack ability of the detection model via fake game, there isn’t sufficient adversarial dynamic beyond generation and evaluation of deepfake speech examples. Future research should develop frameworks that enable dynamic, real-time rivalry game scenarios, allowing for a more thorough exploration of defense mechanisms’ effectiveness and methods like continual reinforcement learning.
@@ VI Future Directions
VII Conclusions
The ADD 2023 challenge aimed to spur innovation and research in detecting and analyzing deepfake speech, attracting 145 teams from 15 countries. This paper presents the challenge’s dataset, provides a technical analysis of top-performing systems and identifies certain limitations and suggests future directions for research in the field of deepfake speech detection, notably in the areas of better coping with unseen attacks, improved interpretability and generalization abilities, real-time processing, multilingual scenarios, and better evaluation metrics. We hope that the release of the dataset and the analysis presented in this paper will inspire further research in the field of deepfake speech detection and contribute to the development of more robust and reliable anti-deepfake technologies.

### v2

@@ 5.3 Track 2: Manipulation region location (RL)
    Acoustic features: Like in the FG-D task, most teams opted to use Wav2Vec 2.0 for acoustic feature extraction, with the notable exception of Teams C02 [43] and C03 [44], who used spectrograms as acoustic feature representation. This popularity of Wav2Vec 2.0 further evidences the reliability of deep embedding features in anti-deepfake systems. In addition, it is worth noting that Team C05 [46] uses both raw waveform as well as wav2vec deep embedding features, an approach that may be more effective in allowing the backend model to both capture the information contained in the raw waveform as well as the deep embedding features, and thus may be more effective in capturing the diverse acoustic conditions of deepfake audio in the wild.
@@
    Back-end classifiers: Given the sequential nature of the RL task, where a real/fake prediction must be given for each segment, and notably due to the fact that audio waveforms are temporal in nature, the use of LSTM-based recurrent models to capture the temporal sequential informations is a popular approach among the top-performing teams, with the notable exception of Team C05 [46], who used an AASIST-based model. Judging by the ranking, however, it is evident that LSTM-based recurrent models may still be more effective and better suited for the recurrent nature of the RL task, as Team C05, who used a fusion of AASIST-based model and FC layer classifier, is the only team in the top-5 that did not use an LSTM-based model, and the performance is slightly lower than that of the other teams.
  4. 4.
    Model fusion: The use of model fusion is not as widespread in Track 2 compared to other tracks, with only C05 opting to use a weighted segment-wise average of the predictions from their models.
5.4 Track 3: Deepfake algorithm recognition (AR)
@@
    Data augmentation: The augmentation of training data to introduce more acoustic conditions, especially with noise and reverberation, is widespread among the top-performing teams. Most teams opted to use MUSAN for noise augmentation, and a significant number of teams also opted to add reverberation to their training data to improve the robustness of their models. In addition, the top-performing teams also opted to use other methods to further augment their training data, including CutMix [58] and random sampling, time stretching, and time masking. This is likely due to the fact that the AR task is concerned with the recognition of deepfake algorithms, and thus the use of different data augmentation methods may be more effective in capturing the diverse acoustic conditions of deepfake audio in the wild.
@@
    Acoustic features: Compared to Tracks 1.2 and 2, participants of Track 3 are more diverse in their choice of acoustic features. While a number of teams used Wav2Vec 2.0, a significant number of teams also opted to use log-Mel spectrogram, STFT spectrogram, and even HuBERT-based features. This is likely due to the fact that the AR task is concerned with the recognition of deepfake algorithms, beyond the binary real/fake classification, and thus a more diverse set of acoustic features may be more suitable, without a clear-cut winner. This also signals the prospect of additional research into the development of acoustic features for anti-deepfake systems.
@@
    Back-end classifiers: ResNet-based [59] models (including Res2Net [60]) are popular among the top-performing teams, with the notable exception of Team D03 [49], who used a single ECAPA-TDNN-based model [61]. The use of ECAPA-TDNN is likely due to the fact that ECAPA-TDNN is a popular model for speaker recognition, and the philosophy of speaker recognition may be applicable to the AR task, which is also a recognition task, albeit of deepfake algorithms instead of speakers. The popularity of ResNet-based models is likely due to the fact that they are relatively simple and efficient, and thus may be more suitable for real-world applications.
@@
    Model fusion: The averaging of scores is still popular, but the use of label fusion and feature fusion (in the form of concatenation) also see its use. This is likely due to the fact that the AR task is concerned with the recognition of deepfake algorithms, and thus the use of different models, each of which may be better suited for a particular aspect of the task, may be more effective than a single-model approach. The use of manifold-based multi-model fusion by Team D05 [51] is particularly interesting, as it allows for the pooling of information from multiple models to make a more informed decision, and thus may be more effective in capturing the diverse acoustic conditions of deepfake audio in the wild.
@@
    OSR methods: Given the nature of the AR task and its practicality in real-world applications, as well as the presence of an unknown class, participating teams are required to use an out-of-distribution detection method to detect unknown classes. Thresholding remains the most popular method, to decent success. Other methods, including OpenMax [62] and manifold-based multi-model fusion, are also used. These methods are effective in detecting unknown classes, likely because they allow for the evaluation of prediction confidence and the pooling of information from multiple models, respectively, in order to make a more informed decision.
6 Future Directions
The ADD 2023 challenge has provided a platform for researchers to develop new technologies to combat deepfake audio; However, there are several limitations that should be addressed in future research efforts. We identify these limitations and suggest future directions for research in the field of deepfake audio detection and analysis.
  • Coping with unseen deepfake technologies and adversarial attacks: The rapid development of audio deepfake generation and adversarial attack technologies, like VALL-E, GPT-4o, VISinger and DiffSinger, etc., brings critical challenges to current existing detection methods. In response to these challenges, deepfake audio generation and detection tasks are viewed as a rivalry game for participants in the ADD competitions. Despite partly improving the anti-attack ability of the detection model via fake game, there isn’t sufficient adversarial dynamic beyond generation and evaluation of deepfake speech examples. Future research should develop frameworks that enable dynamic, real-time rivalry game scenarios, allowing for a more thorough exploration of defense mechanisms’ effectiveness and methods like continual reinforcement learning.
@@ 6 Future Directions
7 Conclusions
The ADD 2023 challenge aimed to spur innovation and research in detecting and analyzing deepfake speech, attracting 145 teams from 15 countries. This paper presents the challenge’s dataset and provides a technical analysis of top-performing systems. Our analysis not only identifies key strengths in the approaches used but also reveals certain limitations that need to be addressed to enhance the robustness of deepfake detection technologies, notably in the areas of better coping with unseen attacks, improved interpretability and generalization abilities, real-time processing, multilingual scenarios, and better evaluation metrics. We hope that the release of the dataset and the analysis presented in this paper will inspire further research in the field of deepfake speech detection and contribute to the development of more robust and reliable anti-deepfake technologies.